OpenStudy (anonymous):

Multiply and simplify: (5y - 2)2

OpenStudy (anonymous):

just distribute

OpenStudy (unklerhaukus):

\[(5y-2)^2\]?

OpenStudy (anonymous):

A.25y2 - 4 B.25y2 + 4 C.25y2 - 10y - 4 D.25y2 - 20y + 4

OpenStudy (unklerhaukus):

\[(5y-2)^2=(5y-2)\times(5y-2)\]\[\qquad\qquad=5y\times(5y-2)-2\times(5y-2)\]\[\qquad\qquad=(5y\times5y)+(5y\times-2)+(-2\times5y)+(-2\times-2)\]

OpenStudy (anonymous):

\[(a+b)^2 = a^2 - 2ab + b^2\] Substitue a and b in the equation, and you get your answer :D

OpenStudy (anonymous):

STILL NEED HELP

OpenStudy (anonymous):

it is 25y2-20y+4

OpenStudy (anonymous):

ie option d

OpenStudy (anonymous):

Where are you stuck ?

OpenStudy (anonymous):

IS THE ANSWER A. B. C. D.????

OpenStudy (anonymous):

yes D is correct

OpenStudy (unklerhaukus):